Problem 5: The greatest possible acceleration or deceleration that a train may have is a and its maximum speed is v. Find the minimum time in which the train can get from one station to the next if the total distance between the stations is s. Assume the train starts from the first station at rest and stops when it reaches the next station.
Answer #1

The acceleration of the train is not uniform, therefore, for t1 time train moved with 'a' acceleration and t 2 time it moved with uniform acceleration, here it is zero and t3 time it moved with deceleration 'a ' that is total time is given as

tmin= t1 + t2 + t3

Then, with 'v' as maximum speed

Ll tmin

t_{min}= 2\frac{v}{a} .............(2)

We know,

v= at

Therefore,

2 min

where S is the distance between the two stations   

Eliminating v from eqn.(3) using eqn.(2), we get

atmin

Therefore ,

4S mat

\mathbf{t_{min}=2\sqrt{\frac{S}{a}} }